You are here: Foswiki>PartnerObserving Web>ItalyQueue2013>ObsInaf20130404 (07 Apr 2013, EleanoraSani)Edit Attach

INAF queue observing, 2013 April 3 - April 4

Observer: E. Sani, M. Faccini, A. Bongiorno
Telescope Operator: D. Gonzales Huerta
Support Astronomer: M. Edwards
Instrument: LBC

Summary

At the beginning of the twilight we have sky with cirrus on the horizon south-west, with 43% humidity and low wind (8-10 m/s).
Good seeing, 1.1, at the beginning of the night operations.
The seeing gets worse during the first scientific program, with an average value of 1.6 in both channel. We thus switch to MODS
Communication problems with M2. The TO needs to calls Doug Miller. After 25 min of efforts by Doug and Runa Briguglio, we are notified that there is no hope to fix the problem in a reasonable time. We thus decide to switch back to LBC.
The seeing gets better after the second switch and stabilizes around 1.0, decreasing till 0.6.
During the Brocato observation there was elongation in the blue channel. Some dofpia reduce the problem for the B images

Data Summary

Project_RUN Objects Filter ExpTime Notes
Giallongo_27E CLJ1120 g-Sloan, r-Sloan 0.5h Bad seeing. To redo
Grazian_30A GOODS-North U-spec, r-Sloan 3h Seeing 1.0 or better
Brocato_DDT Hercules B-Bessel, r-Sloan - U-special, r-Sloan 17+12 min - 8 min The first Br dataset has elongated images in the blue channel; we re-started it within the specs COMPLETE

Details

01:03 The TO opens the dome.

02:08 Start with flats calibrations
  • V from the previous night was saturated. We redo SV
  • B from the previous night was too low. We redo Br

02:09 Run SkyFlatTest_SV_r836d27_pa0_5.ob with time scale =0.5,3

02:13 Run SkyFlat_SV_r836d27_pa0_5.ob with time scale =5

02:18 Run SkyFlat_SV_r836d27_pa180_5.ob with time scale =12
  • After three frames we stop because the counts are too low
  • B from the previous night was too low. We redo Br

02:23 Run SkyFlat_Br_r836d27_pa0_5.ob with time scale =12

02:33 We slew to Giallongo_27E CLJ1120
  • dofpia, /first, /X2 (sky still too bright, stars not visible)

02:41 dofpia, /first not converging

02:46 DIM seeing 2.81. TO cleaning the optics

02:51 dofpia, /first, /X2 not converging

02:57 dofpia, /first, /X2 not converging

03:00 we change field to try to focus. We slew to the photometric standard SA98 airmass=1.3

03:04 dofpia, /first, /X2

03:09 blue channel jumped

03:10 red channel converged but blue channel is producing jellyfish shapes
  • TO cleaning the blue side

03:15 dofpia converged
  • we use this star just to focus

03:20 we slew to Giallongo_27E CLJ1120
  • dofpia, /X2 converged!!
  • Run CLJ1120+43_POINTING.ob
  • lbcrangebal B=033247 R=033242

03:38 Run CLJ1120+43_GR_c.ob
  • The average seeing is ~1.54 in the blue channel and ~1.6 in the red channel.

04:22 We ask the TO to switch to MODS.
  • this is due to the bad and extremely variable seeing that does not allow to obtain images within the specs.

04:38 Problem of communication with the M2.
  • The TO calls Doug

05:08 Dug state that there is no hope any time soon to have the problem fixed. We have to switch back to LBC

05:10 We ask the TO to switch to LBC

05:13 We are authorized to use LBC

05:14 we slew to grazian_GOODSN_UR_21
  • dofpia, /X2
  • Run grazian_GOODSN_pointing.ob
  • lbcrangebal B=052329 R=052326

  • ERROR: left and right preset cancelled!

05:23 rerun grazian_GOODSN_pointing.ob
  • lbcrangebal B=052807 R=052803

05:30 Run grazian_GOODSN_UR_21.ob
  • The average seeing is ~1.1 in both channels

06:10 dofpia, /X2 on the scientific field of Grazian

06:16 Run grazian_GOODSN_UR_22.ob

06:55 the OB got stacked during the last frame
  • LOG ERROR: CCD controller error (the problem seems to be with the tracker)
  • last image not taken

07:10 M. Edwards runs grazian_pointing.ob to be sure everything is working

07:11 M. Edwards RUNS grazian_GOODSN_UR_23.ob and STOP it to check if the telescope is guiding
  • everything seems to work now

07:12 dofpia, /X2 on the scientific field of Grazian

07:14 RUN grazian_GOODSN_UR_23.ob
  • The seeing at the beginning is ~0.8 in the red channel and ~1 in the blue one
  • The seeing at the end is ~1 in both channels

07:54 dofpia, /X2 on the scientific field of Grazian

07:59 RUN grazian_GOODSN_UR_24.ob
  • The seeing at the beginning is ~1 in the red channel and 1.1 in the blue one
  • The seeing at the end is 1 in both channels

08:45 dofpia, /X2 on the scientific field of Grazian

08:51 RUN grazian_GOODSN_UR_25.ob
  • The seeing at the beginning is ~0.8 in the red channel and 1.0 in the blue one
  • The seeing at the end is 0.8 in both channels

09:31 dofpia, /X2 on the scientific field of Grazian

09:35 RUN grazian_GOODSN_UR_26.ob
  • The seeing at the beginning is ~0.8 in both channels
  • The seeing at the end is 0.9 in both channels

10:17 Slewing to Brocato_DDT Hercules
  • Run hercules1_pointing.ob
  • dofpia, /X2
  • lbcrangebal B=102734 R=102729

10:30 Run hercules1_Br.ob
  • The seeing is 0.8 in the red channel and 0.9 in the blue one at the beginning of the integration.
  • Slightly elongate images in the blue channel. The issue could be related to a quick increase of the temperature; thus the mirrors are not able to follow such T gradient.

11:02 Run hercules2_pointing.ob
  • dofpia, /X2 on scientific field Brocato
  • lbcrangebal B=110853 R=110848

11:13 Run hercules2_Br.ob
  • In both channels, the seeing is 0.8 at the beginning of the integration and 0.9 at the end.
  • No elongation

11:29 Run hercules1_Br.ob
  • since the last OB has no elongated feature, we re-run the first filter settings.
  • In both channels, the seeing is 0.9
  • due to elongation in blue channel we run dofpia, /X2 on scientific field Brocato

11:40 Run hercules1_Br.ob
  • The seeing is ~1.0 in the red channel and ~1.1 in the blue one at the beginning of the integration.

12:00 We stop hercules1_Br.ob after 24/34 frames due to twilight

12:04 Start with flats calibrations
  • B from the previous night was too low. We redo Br

12:05 we run some test skyFlat

12:25 Run SkyFlat_gI_r1731d_pa0_5.ob with time scale = 8

12:29 Run SkyFlat_gI_r1731d_pa180_5.ob with time scale = 3

12:37 Run SkyFlat_Br_r1731d_pa180_5.ob with time scale = 5

12:45 The TO close the dome
  • Run 25Bians_bino.ob

-- MarcoFaccini - 04 Apr 2013
Topic revision: r8 - 07 Apr 2013, EleanoraSani
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback